# Collaborative Face Experts Fusion in Video Generation: Boosting Identity Consistency Across Large Face Poses

### v3

@@ 1 Introduction
Recent Diffusion Transformer (DiT)-based image-to-video models such as CogVideoX [yang2024cogvideox] and Wan Model [wan2025wan] achieve content consistency by injecting first-frame VAE features into the DiT backbone. To better preserve character identity, human-centric approaches like ConsisID [yuan2025consisid] further inject a frequency-aware facial representation into DiT. However, both paradigms share critical limitations: they rely on a single face feature extracted from the reference frame and inject the same static face feature uniformly across all timesteps and DiT layers, thereby limiting the model’s ability to capture discriminative facial details. This leads to unstable identity preservation under large face poses, often resulting in distorted character appearances. Furthermore, existing human-centric datasets [wang2024humanvid, li2025openhumanvid] are ill-suited for learning pose-robust identity preservation. They often feature faces of inconsistent quality, lack sufficient dynamic large-angle examples, and seldom ensure cross-frame identity consistency or sufficient facial visibility, hindering the learning of pose-invariant identity representations.
@@
Complementing our architectural innovation, we introduce LaFID-180K (Large Face Pose for Identity Preservation Dataset), a curated large-scale collection of video clips annotated with 3D face poses. This dataset specifically targets the scarcity of high-quality, identity-consistent videos under large face poses. LaFID-180K is built through a three-stage pipeline that directly addresses key data limitations: Face Constraints ensure clear facial visibility and diverse large-pose coverage; Identity Consistency enforces temporal identity coherence via feature clustering; and Speech Disambiguation aligns video captions with lip movements by verifying speaking status, ensuring accurate annotations under pose-induced ambiguity. We evaluate our method on two distinct benchmarks: the existing VIP-Test [vip-200k] for general full-body motion, and our newly introduced LaFID-Bench, which addresses the underexplored setting of large face poses.
@@ 2.1 Identity-Preserving Text-to-Video Generation
In the identity-preserving text-to-video (IPT2V) task, early methods used time-consuming fine-tuning for facial identity preservation [chefer2024still, wu2024motionbooth, wei2024dreamvideo, zhong2025concat, lei2024comprehensive, ma2025controllable], driving tuning-free approaches to dominance. Pioneering works like ID-Animator [he2024idanimator] relied on static image-model features, harming dynamic identity consistency. ConsisID [yuan2025consisid] injected facial frequency signals but suffered from 2D pose mismatches; FantasyID [zhang2025fantasyid] added 3D priors but lacked adaptive feature selection. Recent methods like VACE [jiang2025vace] and SkyReels-A2 [fei2025skyreels] used reference branches but failed to dynamically prioritize cues during large motions. These show I2V needs adaptive fusion of face features aligned with motion, especially large face motions. Thus, we introduce Collaborative Face Experts Fusion (CoFE), a DiT layer-adaptive gating mechanism that dynamically fuses multiple features to overcome static fusion and enhance large-pose robustness.
@@ 2.2 Human-Centric Image-to-Video Generation
Image-to-video (I2V) generation [guo2024i2v, bharadhwaj2024gen2act, xu2025freevis] struggles to control content reliably with text alone, so recent methods use first-frame conditioning for better controllability. AnimateAnything [dai2023animateanything] and PixelDance [zeng2024make] modify U-Nets by concatenating first-frame latent features with input noise to anchor spatial features for temporal consistency. Dynamicrafter [xing2024dynamicrafter] and Moonshot [zhang2024moonshot] enhance conditioning with image cross-attention layers, injecting stronger visual cues into diffusion. Recent methods like Wan2.1, Wan2.2 [wan2025wan], and VACE [jiang2025vace] support I2V with strong performance but struggle to preserve identity across large face poses. Moreover, IPRO [shen2025identity] introduces reinforcement learning to enhance training for facial details. Generally, identity robustness under large face pose variations remains largely unaddressed in current I2V frameworks.
@@ 2.3 Human-Centric Video Datasets
Large-scale human-centric video datasets [sadoughi2015msp, chen2024panda, bain2021frozen] are foundational for human video generation, yet most prioritize body motion or general scenes over facial identity under large poses. For example, TikTok [jafarian2021learning] focuses on social dance; UCF101 [soomro2012ucf101] and NTU RGB+D [shahroudy2016ntu] on action recognition; HumanVid [wang2024humanvid] on synthetic-real blends; and OpenHumanVid [li2025openhumanvid] provides high-resolution videos but lacks large face pose diversity, facial quality filtering, and fine-grained 3D face pose annotations, limiting its suitability for pose-robust identity learning. To address it, we introduce LaFID-180K, a dataset constructed by combining OpenHumanVid [li2025openhumanvid] with carefully selected in-house videos, and curated through a unified pipeline to emphasize diverse large face poses with 3D face pose annotations.
@@ 3 Methodology: CoFE
Current image-to-video models [yang2024cogvideox, hacohen2024ltx, wan2025wan, jiang2025vace] inject identity using a single-source feature extracted from the first frame, which is then applied statically across all timesteps and network layers, typically via uniform cross-attention. This design suffers from two fundamental limitations. First, the single feature conflates identity, semantics, and details into one representation, lacking specialization. Second, its static injection cannot adapt to pose-induced appearance changes or layer-specific attention demands, leading to severe identity drift under large pose variations. To overcome this, we propose Collaborative Face Experts Fusion (CoFE), a layer-adaptive fusion module that dynamically integrates three specialized facial representations into cross-attention: pose-robust identity features, semantic attributes (e.g., hairstyle, expression), and pixel-level details (e.g., color gradients). An overview of our pipeline is shown in Fig. 2.
@@
Motivation from Cross-Attention Patterns. To understand how different facial representations influence identity preservation, we analyze cross-attention patterns in the generator. As shown in Figure 3, each representation induces a distinct attention pattern: the Identity Expert (ArcFace [deng2019arcface]) focuses on pose-invariant landmarks such as the eyes, nose, and mouth to capture structural identity; the Semantic Expert (CLIP [radford2021learning]) attends to high-level attributes including hairstyle and expression; and the Detail Expert (DINOv2 [oquab2023dinov2]) highlights fine-grained textures and structural edges. These complementary roles indicate that robust identity preservation under large pose variations requires dynamic fusion of all three signals rather than reliance on a single static feature.
@@
Koala-36M [wang2025koala] Open 36M - 720P
OpenHumanVid [li2025openhumanvid] Human 13.2M Skeleton pose 720P
@@
Collaborative Face Experts Fusion. Our goal is to fuse hierarchical face features adaptive to distinct DiT blocks, thereby achieving complementary synergies that enhance identity persistence, semantic coherence, and pixel-level fidelity. To achieve this, we first employ a face preprocessor comprising a face detector [deng2020retinaface] and a face parsing model [geitgey2019face] to obtain the cropped and aligned face image from the input image, removing background to focus on the face region. These preprocessed faces are then processed by identity, semantic, and detail encoders to generate identity-sensitive features, pixel-level detail features, and high-level semantic features, denoted as $\mathbf{f}_{\text{id}},\mathbf{f}_{\text{sem}},\mathbf{f}_{\text{det}}$, respectively.
@@ 4.1 Overview
Existing benchmarks such as VIP-Test [vip-200k] focus on full-body motion but lack sufficient variation in large face poses, limiting their ability to evaluate identity preservation under challenging pose conditions. To address this gap, we introduce LaFID-Bench, a standardized evaluation protocol derived from LaFID-180K, comprising fifty held-out identities (disjoint from training) with four diverse prompts per identity, all featuring large face poses that span extreme yaw and pitch ranges. For generalization assessment, we additionally construct a VIP-Test subset using SDXL-generated [podell2023sdxl] reference images (fifty identities, four prompts each). Together, these two test sets enable the first holistic evaluation of image-to-video models across both standard full-body animation and the challenging regime of identity preservation under large face poses.
@@ 4.2 Data Processing Pipeline
CogVideoX-5B [yang2024cogvideox] 0.629 0.636 29.331 104.214 0.499 0.598 0.976 0.986
LTX-2B [hacohen2024ltx] 0.438 0.440 28.800 151.773 0.513 0.581 0.980 0.992
ConsisID [yuan2025consisid] 0.567 0.573 31.184 123.967 0.560 0.670 0.979 0.990
Wan2.1-I2V-1.3B [wan2025wan] 0.654 0.656 29.200 72.322 0.553 0.660 0.989 0.993
@@
CogVideoX-5B [yang2024cogvideox] 0.462 0.474 28.947 163.001 0.512 0.585 0.926 0.948
LTX-2B [hacohen2024ltx] 0.301 0.323 28.403 203.516 0.510 0.586 0.944 0.975
ConsisID [yuan2025consisid] 0.269 0.289 29.601 206.924 0.566 0.638 0.946 0.976
Wan2.1-I2V-1.3B [wan2025wan] 0.471 0.480 29.825 129.116 0.554 0.659 0.940 0.967
@@
Face Constraints Filtering. To support identity preservation under diverse face poses, especially large face motions, we curate high-quality samples for LaFID via multi-dimensional face constraints filtering, enforcing strict criteria on face continuity, visibility, and pose diversity. For efficiency, we sample one frame every three and detect face bounding boxes and poses using ArcFace [deng2020retinaface]. The core constraints are: i) Face count. Focused on single person generation, we filter videos where any selected frame contains no faces or multiple faces. This ensures a single, continuous face throughout the video clip, foundational for consistent identity learning. ii) Face proportion. Video clips with low face bounding box proportion (i.e., less than 10%) are filtered out to emphasize the preservation of facial details. iii) Face pose diversity. To ensure preservation of identity across all face poses, we retain videos with diverse poses. We analyze the face poses of selected frames for each video clip, selecting videos with large face pose variations.
Identity Consistency Analysis Identity consistency is essential for human-centric videos, reflecting the stability of facial representations across frames. To ensure high-quality data curation, we analyze identity coherence using ArcFace features [deng2019arcface]. i) Identity Similarity. We compute pairwise cosine similarities between frame embeddings $\{\mathbf{d}_{i}\}_{i=1}^{N}$:
@@
and use the average off-diagonal similarity $\overline{s}$ as a measure of intra-clip identity coherence. Clips with $\overline{s}\leq 0.6$ are discarded. ii) Identity Clustering. To ensure disjoint identity splits for fair train/test evaluation, we cluster videos by identity using FAISS [douze2024faiss]. An IVF-PQ index [xiao2022distill] enables efficient similarity search in the ArcFace embedding space via product quantization. A two-stage hierarchical strategy is applied: a high threshold first forms tight identity clusters from strongly correlated samples; a lower threshold then assigns ambiguous samples to recover weakly correlated but same-identity instances.
Speech Disambiguation. LaFID is mainly curated from OpenHumanVid [li2025openhumanvid], where most videos show people speaking with continuous lip movements but captions rarely indicate actual speech or lip motion. Without explicit speech semantics, identity-preserving generators often wrongly link lip motion to text, producing videos with incessant lip movement even when no speech motion is contained in the input caption. To align lip motion with caption, we disentangle speaking status in two steps: whisper [radford2023robust] detects audible speech and its alignment with the primary subject’s lip movements; if the caption lacks explicit description of speaking, an LLM [bai2023qwen, touvron2023llama] appends a context-aware sentence (e.g., “The man is talking.”) to signal active speech. This ensures genuine speaking behavior is reflected in text-video alignment, enabling controllable and identity-consistent generation.
@@ 5.1.1 Training strategies
We integrate the proposed CoFE into the Wan2.1-I2V-1.3B and Wan2.2-TI2V-5B models [wan2025wan]. To balance performance and efficiency, we inject CoFE into each even-numbered DiT block, in line with VACE [jiang2025vace]. This design enhances model capabilities while incurring a modest parameter increase of merely 18.5%. We employ aspect ratio bucketing for multi-resolution and multi-frame training, with a minimum video resolution of 480p and a frame count ranging from 49 to 129. To address the “copy-paste” shortcut, we incorporate face features from non-training frames into the input for video clips with additional frames. We use the Adam optimizer with a learning rate of 1e-5 while keeping other modules frozen, and train on the LaFID-180K dataset for 4,000 iterations with a batch size of 256.
@@ 5.1.2 Evaluation Protocols
We evaluate our method on LaFID-Bench and VIP-Test [vip-200k], adopting the evaluation protocol of ConsisID [yuan2025consisid] for metric selection. Specifically, we assess identity preservation using FaceSim-Arc [deng2019arcface] and FaceSim-Cur [huang2020curricularface], text relevance via CLIP-Score [hessel2021clipscore], facial fidelity through FID, and overall video quality with VBench [huang2024vbench]. The latter includes four key sub-metrics: Aesthetic Quality (AQ), Imaging Quality (IQ), Temporal Flickering (TF), and Motion Smoothness (MS).
@@ 5.2 Benchmark Results
We compare our method with open-source state-of-the-art approaches [yang2024cogvideox, hacohen2024ltx, yuan2025consisid, wan2025wan] on LaFID-Bench and VIP-Test [vip-200k] to demonstrate the superiority of our CoFE module in preserving identity under large face poses.
@@ 6 Conclusion
In this paper, we address facial identity preservation under large pose variations in image-to-video generation through two key contributions. First, we introduce the Collaborative Face Experts Fusion (CoFE) module, which dynamically fuses identity-sensitive, semantic, and detail-oriented face features through a layer-adaptive cross-attention gating mechanism. Second, we curate LaFID-180K through a three-stage pipeline: facial constraints filtering for large-pose diversity, identity consistency analysis for cross-frame identity stability, and speech disambiguation to align captions with speaking status, effectively reducing lip motion and text ambiguity. To support rigorous evaluation in the challenging regime of large face poses, we introduce LaFID-Bench, a new benchmark specifically designed for this scenario. Experiments show that CoFE achieves state-of-the-art performance on both the general-purpose VIP-Test [vip-200k] and our specialized LaFID-Bench, demonstrating consistent improvements in identity preservation and overall video quality across diverse generation settings.

### v4

@@ 1 Introduction
Recent Diffusion Transformer (DiT)-based image-to-video models such as CogVideoX [41] and Wan Model [33] achieve content consistency by injecting first-frame VAE features into the DiT backbone. To better preserve character identity, human-centric approaches like ConsisID [42] further inject a frequency-aware facial representation into DiT. However, both paradigms share critical limitations: they rely on a single face feature extracted from the reference frame and inject the same static face feature uniformly across all timesteps and DiT layers, thereby limiting the model’s ability to capture discriminative facial details. This leads to unstable identity preservation under large face poses, often resulting in distorted character appearances. Furthermore, existing human-centric datasets [35, 21] are ill-suited for learning pose-robust identity preservation. They often feature faces of inconsistent quality, lack sufficient dynamic large-angle examples, and seldom ensure cross-frame identity consistency or sufficient facial visibility, hindering the learning of pose-invariant identity representations.
@@
Complementing our architectural innovation, we introduce LaFID-180K (Large Face Pose for Identity Preservation Dataset), a curated large-scale collection of video clips annotated with 3D face poses. This dataset specifically targets the scarcity of high-quality, identity-consistent videos under large face poses. LaFID-180K is built through a three-stage pipeline that directly addresses key data limitations: Face Constraints ensure clear facial visibility and diverse large-pose coverage; Identity Consistency enforces temporal identity coherence via feature clustering; and Speech Disambiguation aligns video captions with lip movements by verifying speaking status, ensuring accurate annotations under pose-induced ambiguity. We evaluate our method on two distinct benchmarks: the existing VIP-Test [23] for general full-body motion, and our newly introduced LaFID-Bench, which addresses the underexplored setting of large face poses.
@@ 2.1 Identity-Preserving Text-to-Video Generation
In the identity-preserving text-to-video (IPT2V) task, early methods used time-consuming fine-tuning for facial identity preservation [4, 37, 36, 46, 20, 22], driving tuning-free approaches to dominance. Pioneering works like ID-Animator [14] relied on static image-model features, harming dynamic identity consistency. ConsisID [42] injected facial frequency signals but suffered from 2D pose mismatches; FantasyID [45] added 3D priors but lacked adaptive feature selection. Recent methods like VACE [19] and SkyReels-A2 [10] used reference branches but failed to dynamically prioritize cues during large motions. These show I2V needs adaptive fusion of face features aligned with motion, especially large face motions. Thus, we introduce Collaborative Face Experts Fusion (CoFE), a DiT layer-adaptive gating mechanism that dynamically fuses multiple features to overcome static fusion and enhance large-pose robustness.
@@ 2.2 Human-Centric Image-to-Video Generation
Image-to-video (I2V) generation [12, 3, 40] struggles to control content reliably with text alone, so recent methods use first-frame conditioning for better controllability. AnimateAnything [6] and PixelDance [43] modify U-Nets by concatenating first-frame latent features with input noise to anchor spatial features for temporal consistency. Dynamicrafter [39] and Moonshot [44] enhance conditioning with image cross-attention layers, injecting stronger visual cues into diffusion. Recent methods like Wan2.1, Wan2.2 [33], and VACE [19] support I2V with strong performance but struggle to preserve identity across large face poses. Moreover, IPRO [30] introduces reinforcement learning to enhance training for facial details. Generally, identity robustness under large face pose variations remains largely unaddressed in current I2V frameworks.
@@ 2.3 Human-Centric Video Datasets
Large-scale human-centric video datasets [28, 5, 2] are foundational for human video generation, yet most prioritize body motion or general scenes over facial identity under large poses. For example, TikTok [18] focuses on social dance; UCF101 [31] and NTU RGB+D [29] on action recognition; HumanVid [35] on synthetic-real blends; and OpenHumanVid [21] provides high-resolution videos but lacks large face pose diversity, facial quality filtering, and fine-grained 3D face pose annotations, limiting its suitability for pose-robust identity learning. To address it, we introduce LaFID-180K, a dataset constructed by combining OpenHumanVid [21] with carefully selected in-house videos, and curated through a unified pipeline to emphasize diverse large face poses with 3D face pose annotations.
@@ 3 Methodology: CoFE
Current image-to-video models [41, 13, 33, 19] inject identity using a single-source feature extracted from the first frame, which is then applied statically across all timesteps and network layers, typically via uniform cross-attention. This design suffers from two fundamental limitations. First, the single feature conflates identity, semantics, and details into one representation, lacking specialization. Second, its static injection cannot adapt to pose-induced appearance changes or layer-specific attention demands, leading to severe identity drift under large pose variations. To overcome this, we propose Collaborative Face Experts Fusion (CoFE), a layer-adaptive fusion module that dynamically integrates three specialized facial representations into cross-attention: pose-robust identity features, semantic attributes (e.g., hairstyle, expression), and pixel-level details (e.g., color gradients). An overview of our pipeline is shown in Fig. 2.
@@
Motivation from Cross-Attention Patterns. To understand how different facial representations influence identity preservation, we analyze cross-attention patterns in the generator. As shown in Figure 3, each representation induces a distinct attention pattern: the Identity Expert (ArcFace [7]) focuses on pose-invariant landmarks such as the eyes, nose, and mouth to capture structural identity; the Semantic Expert (CLIP [26]) attends to high-level attributes including hairstyle and expression; and the Detail Expert (DINOv2 [24]) highlights fine-grained textures and structural edges. These complementary roles indicate that robust identity preservation under large pose variations requires dynamic fusion of all three signals rather than reliance on a single static feature.
@@
Koala-36M [34] Open 36M - 720P
OpenHumanVid [21] Human 13.2M Skeleton pose 720P
@@
Collaborative Face Experts Fusion. Our goal is to fuse hierarchical face features adaptive to distinct DiT blocks, thereby achieving complementary synergies that enhance identity persistence, semantic coherence, and pixel-level fidelity. To achieve this, we first employ a face preprocessor comprising a face detector [8] and a face parsing model [11] to obtain the cropped and aligned face image from the input image, removing background to focus on the face region. These preprocessed faces are then processed by identity, semantic, and detail encoders to generate identity-sensitive features, pixel-level detail features, and high-level semantic features, denoted as $\mathbf{f}_{\text{id}},\mathbf{f}_{\text{sem}},\mathbf{f}_{\text{det}}$, respectively.
@@ 4.1 Overview
Existing benchmarks such as VIP-Test [23] focus on full-body motion but lack sufficient variation in large face poses, limiting their ability to evaluate identity preservation under challenging pose conditions. To address this gap, we introduce LaFID-Bench, a standardized evaluation protocol derived from LaFID-180K, comprising fifty held-out identities (disjoint from training) with four diverse prompts per identity, all featuring large face poses that span extreme yaw and pitch ranges. For generalization assessment, we additionally construct a VIP-Test subset using SDXL-generated [25] reference images (fifty identities, four prompts each). Together, these two test sets enable the first holistic evaluation of image-to-video models across both standard full-body animation and the challenging regime of identity preservation under large face poses.
@@ 4.2 Data Processing Pipeline
CogVideoX-5B [41] 0.629 0.636 29.331 104.214 0.499 0.598 0.976 0.986
LTX-2B [13] 0.438 0.440 28.800 151.773 0.513 0.581 0.980 0.992
ConsisID [42] 0.567 0.573 31.184 123.967 0.560 0.670 0.979 0.990
Wan2.1-I2V-1.3B [33] 0.654 0.656 29.200 72.322 0.553 0.660 0.989 0.993
@@
CogVideoX-5B [41] 0.462 0.474 28.947 163.001 0.512 0.585 0.926 0.948
LTX-2B [13] 0.301 0.323 28.403 203.516 0.510 0.586 0.944 0.975
ConsisID [42] 0.269 0.289 29.601 206.924 0.566 0.638 0.946 0.976
Wan2.1-I2V-1.3B [33] 0.471 0.480 29.825 129.116 0.554 0.659 0.940 0.967
@@
Face Constraints Filtering. To support identity preservation under diverse face poses, especially large face motions, we curate high-quality samples for LaFID via multi-dimensional face constraints filtering, enforcing strict criteria on face continuity, visibility, and pose diversity. For efficiency, we sample one frame every three and detect face bounding boxes and poses using ArcFace [8]. The core constraints are: i) Face count. Focused on single person generation, we filter videos where any selected frame contains no faces or multiple faces. This ensures a single, continuous face throughout the video clip, foundational for consistent identity learning. ii) Face proportion. Video clips with low face bounding box proportion (i.e., less than 10%) are filtered out to emphasize the preservation of facial details. iii) Face pose diversity. To ensure preservation of identity across all face poses, we retain videos with diverse poses. We analyze the face poses of selected frames for each video clip, selecting videos with large face pose variations.
Identity Consistency Analysis Identity consistency is essential for human-centric videos, reflecting the stability of facial representations across frames. To ensure high-quality data curation, we analyze identity coherence using ArcFace features [7]. i) Identity Similarity. We compute pairwise cosine similarities between frame embeddings $\{\mathbf{d}_{i}\}_{i=1}^{N}$:
@@
and use the average off-diagonal similarity $\overline{s}$ as a measure of intra-clip identity coherence. Clips with $\overline{s}\leq 0.6$ are discarded. ii) Identity Clustering. To ensure disjoint identity splits for fair train/test evaluation, we cluster videos by identity using FAISS [9]. An IVF-PQ index [38] enables efficient similarity search in the ArcFace embedding space via product quantization. A two-stage hierarchical strategy is applied: a high threshold first forms tight identity clusters from strongly correlated samples; a lower threshold then assigns ambiguous samples to recover weakly correlated but same-identity instances.
Speech Disambiguation. LaFID is mainly curated from OpenHumanVid [21], where most videos show people speaking with continuous lip movements but captions rarely indicate actual speech or lip motion. Without explicit speech semantics, identity-preserving generators often wrongly link lip motion to text, producing videos with incessant lip movement even when no speech motion is contained in the input caption. To align lip motion with caption, we disentangle speaking status in two steps: whisper [27] detects audible speech and its alignment with the primary subject’s lip movements; if the caption lacks explicit description of speaking, an LLM [1, 32] appends a context-aware sentence (e.g., “The man is talking.”) to signal active speech. This ensures genuine speaking behavior is reflected in text-video alignment, enabling controllable and identity-consistent generation.
@@ 5.1.1 Training strategies
We integrate the proposed CoFE into the Wan2.1-I2V-1.3B and Wan2.2-TI2V-5B models [33]. To balance performance and efficiency, we inject CoFE into each even-numbered DiT block, in line with VACE [19]. This design enhances model capabilities while incurring a modest parameter increase of merely 18.5%. We employ aspect ratio bucketing for multi-resolution and multi-frame training, with a minimum video resolution of 480p and a frame count ranging from 49 to 129. To address the “copy-paste” shortcut, we incorporate face features from non-training frames into the input for video clips with additional frames. We use the Adam optimizer with a learning rate of 1e-5 while keeping other modules frozen, and train on the LaFID-180K dataset for 4,000 iterations with a batch size of 256.
@@ 5.1.2 Evaluation Protocols
We evaluate our method on LaFID-Bench and VIP-Test [23], adopting the evaluation protocol of ConsisID [42] for metric selection. Specifically, we assess identity preservation using FaceSim-Arc [7] and FaceSim-Cur [16], text relevance via CLIP-Score [15], facial fidelity through FID, and overall video quality with VBench [17]. The latter includes four key sub-metrics: Aesthetic Quality (AQ), Imaging Quality (IQ), Temporal Flickering (TF), and Motion Smoothness (MS).
@@ 5.2 Benchmark Results
We compare our method with open-source state-of-the-art approaches [41, 13, 42, 33] on LaFID-Bench and VIP-Test [23] to demonstrate the superiority of our CoFE module in preserving identity under large face poses.
@@ 6 Conclusion
In this paper, we address facial identity preservation under large pose variations in image-to-video generation through two key contributions. First, we introduce the Collaborative Face Experts Fusion (CoFE) module, which dynamically fuses identity-sensitive, semantic, and detail-oriented face features through a layer-adaptive cross-attention gating mechanism. Second, we curate LaFID-180K through a three-stage pipeline: facial constraints filtering for large-pose diversity, identity consistency analysis for cross-frame identity stability, and speech disambiguation to align captions with speaking status, effectively reducing lip motion and text ambiguity. To support rigorous evaluation in the challenging regime of large face poses, we introduce LaFID-Bench, a new benchmark specifically designed for this scenario. Experiments show that CoFE achieves state-of-the-art performance on both the general-purpose VIP-Test [23] and our specialized LaFID-Bench, demonstrating consistent improvements in identity preservation and overall video quality across diverse generation settings.